The radius of the fifth zone of a zone plate is 2 mm. Considering the zone plate as

a converging lens, calculate its focal length for light of wavelength 4800 Å.


1
Expert's answer
2021-05-04T06:50:16-0400


Answer

Given data in question

"r_5=0.002m"

n=5

"\lambda=4800\times10^{-10}m"

Thus, the focal length of the zone plate,

"f =\frac{ r_n^2}{n\lambda}\\=
\frac{ (0.002)^2}{5\times 4800\times10^{-10} }
\\=166.67m"